# Fine-tuning the ESM2 protein language model to understand the functional impact of missense variants

### v1

@@ Appendix A Appendix
We used ACMG/AMP guidelines (Richards et al., 2015) to classify the variants into putative pathogenicity groups, as described in our previous works (Saadat et al., 2023; Saadat & Fellay, 2024a, b). In summary, we gather all the available evidences for a variant. Table 1 summarizes all the ACMG/AMP criteria that we used.
Table 1: the summary of ACMG/AMP criteria used for variant classification. MAF: minor allele frequency
To calculate the probability of pathogenicity (PoP), we use the Bayesian framework developed by Tavtigian et al. (2018). For a given variant, the PoP is calculated as follow:
@@
Supplementary Figure S1: Number of annotated proteins per feature
Supplementary Figure S2: Number of labeled/predicted proteins and amino acids
Supplementary Figure S3: Distribution of conservation score for labeled/predicted amino acids
Supplementary Figure S4: Distribution of pathogenicity score for labeled/predicted amino acids

### v2

@@ Appendix A Appendix
Probability of pathogenicity (PoP) calculation
We used ACMG/AMP guidelines 4 to classify the variants into putative pathogenicity groups, as described in our previous works 30, saadat-fellay-2024-dna. In summary, we gather all the available evidences for a variant. Figure S4 summarizes all the ACMG/AMP criteria that we used.
Supplementary Figure S4: the summary of ACMG/AMP criteria used for variant classification. MAF: minor allele frequency
To calculate the probability of pathogenicity (PoP), we use the Bayesian framework developed by Tavtigian et al. 31. For a given variant, the PoP is calculated as follow: